OpenStudy (anonymous):

can somebody plz help i will give a medal and a fan :) write three ratios equal to 4/36 A. 1/8 - 2/16- 3/24 B. 4/36 - 4/40 - 4/44 C. 1/9 - 2/18 - 3/27 D.1/9 - 1/18 - 1/27

OpenStudy (anonymous):

4/36 = 4 into 36 is 9. Reason 4 * 9 = 36 4/36 = 1/9 So answer is (C).

OpenStudy (anonymous):

thnx